+ + 9 = 0

k values and general solution
Answer #1

Sol. dzy dy + qy=0 -lis Auxillary equation ou | mitm+9=0 m= -1 + $1092-469)(1) 2(1) 09 m = -1 + ST-36 - - 1 + 135 Gyoma Gener